OpenStudy (caerus):

can anyone help me to solve this bernoulli's equation or substitution..ill give medal

OpenStudy (caerus):

\[(3tanx-2cosy)\sec^2xdx+tanxsinydy=0\]

OpenStudy (eliesaab):

Can you finish it?

OpenStudy (caerus):

ill try

OpenStudy (caerus):

its hard to analyze that thing lol

OpenStudy (caerus):

@SolomonZelman

OpenStudy (holsteremission):

If you're not familiar with the method of exact equations, I think the substitution you need here is \(z=\cos y\), which gives \(\mathrm dz=-\sin y\,\mathrm dy\). This changes the ODE to \[(3\tan x-2z)\sec^2x\,\mathrm dx-\tan x\,\mathrm dz=0\]Another substitution of \(u=\tan x\) gives \(\mathrm du=\sec^2x\,\mathrm dx\), so you now have \[(3u-2z)\,\mathrm du-u\,\mathrm dz=0\iff u\frac{\mathrm dz}{\mathrm du}+2z=3u\]which is linear.